[figure: 9.7 cm]

FINITE-TEMPERATURE QUANTUM TRANSFER-MATRIX SIMULATIONS OF THE FRUSTRATED SPIN 1/2 CHAINS

G. Kamieniarz^{1}, M. Bieliński^{1}, G. Szukowski^{1} H. Szymczak^{2}, M. Baran^{2}

1Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań,

Poland 2Institute of Phisics, Polish Academy of Sciences, Warsaw, Poland

Thermodynamical properties of the one-dimensional S = 1/2 Heisenberg model with nearest and next-nearest neighbour interactions are studied by the numerically exact quantum transfer-matrix method and the results are applied to the quasi-one-dimensional magnetic systems CuGeO_{3} and Pb[Cu(SO_{4})(OH)_{2}]. The Suzuki-Trotter formula is used to obtain a classical system with spin σ = 3/2 and effective interactions between nearest neigbours only [1]. The magnetic susceptibility curve is calculated and compared with experimental results performed in a wide temperature range, revealing the presence of frustration in the model proposed for CuGeO_{3} and Pb[Cu(SO_{4})(OH)_{2}]. For the pure CuGeO_{3}, temperature dependence of the dimerization parameter below the spin-Peierls transition point is also estimated. The technique is extended for the finite-chain segments and applied to the diluted CuGeO_{3} obtained by the neutron beam with three different intensities. The corresponding sizes of segments are also estimated for all the samples.

[1] G. Kamieniarz, M. Bieliński and J.P. Renard, Phys. Rev. B 60, 14521 (1999)

[figure: 13.4 cm]

Subject category :

6. Theory of Magnetism

Presentation mode : poster

Corresponding author : G. Kamieniarz

Address for correspondence : Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

Email address : gjk@amu.edu.pl